i. A fixed mass of gas has a pressure of 1.2 MPa and volume of 0.3 m3 and temperature of 100 oC . The pressure is changed to 400 kPa and the volume changed to 0.06 m3. Determine the new temperature?

 

ii. A mass of gas is stored in a cylinder. The gas is stored at a pressure of 1.2 MPa and the cylinder has a volume of 0.03 m3 and temperature of 100 oC , Find the mass of the gas. The characteristic gas constant for air is 300 Jkg−1K−1.



1
Expert's answer
2021-05-07T08:12:38-0400

i. Using the general gas law we find

p1V1/T1 = p2V2/T2

where

p1 = 1.2 x 106 Pa

V1 = 0.3 m3

p2 = 400 x 103 Pa

V2 = 0.06 m3

T1 = (100 + 273.15) K = 373.15 K

T2 = (p2V2T1) / (p1V1) = (400 x 103 Pa x 0.06 m3 x 373.15 K) / (1.2 x 106 Pa x 0.3 m3) = 24.88 K


ii. pV = nRT

1.2 x 106 Pa x 0.03 m3 = m x 300 Jkg−1K−1 x 373.15 K

m = 0.322 kg